OpenStudy (anonymous):

Is that given that 2/3 , 8/d , n/39 are equilvalent fractions . Find the value of d and the value of n. Anybody help me :) TY

OpenStudy (anonymous):

2/3 = 8/d and cross-products produce 2d = 24. DIvide each side by 2 and d = 12. 2/3 = n/39 and cross-products produce 3n = 78. Divide each side by 3 and n = 26.

OpenStudy (anonymous):

thank You :)